OpenStudy (anonymous):

Find roots of the polynomial P(x) = x^3 + 4x^2 - 4x - 16

OpenStudy (anonymous):

this polynomial is factorable by grouping; can you factor by grouping?

OpenStudy (anonymous):

Not really. Factoring always has confused me

OpenStudy (anonymous):

\[x^3+4x^2-4x-16=x^2(x+4)-1(x+4)\]

OpenStudy (anonymous):

do you see how i did that? or no?

OpenStudy (anonymous):

I really hate to look stupid, but I have no idea what you did to get that.

OpenStudy (anonymous):

if you knew how to do it you wouldn't be here :) i'm here to help you so ask whatever you want ok? look at the first two terms, what factor do they have in common?

OpenStudy (anonymous):

x^2?

OpenStudy (anonymous):

yes! and i took out the x^2 and (x+4) is left that's where x^2(x+4) comes from. now the last two terms have -x - 4 the only thing common there is -1. So the first step factors into \[x^2(x+4)-1(x+4)\]

OpenStudy (anonymous):

now what do those two terms have in common?

OpenStudy (anonymous):

Why do you have the z?

OpenStudy (anonymous):

I don't see a z, you mean \[x\]

OpenStudy (anonymous):

Ha, yeah! What is that?? It's confusing.

OpenStudy (anonymous):

thats an X

OpenStudy (anonymous):

is this better? \[X^2(X+4)-1(X+4)\]

OpenStudy (anonymous):

MUCH

OpenStudy (anonymous):

ok now what's in common?

OpenStudy (anonymous):

x+4

OpenStudy (anonymous):

i made a mistake because i copied the problem down wrong its minor but it should be \[X^2(X+4)-4(X+4)\]

OpenStudy (anonymous):

left a 4 off... but yes (X+4) is the common factor. When I've tutored this is usually the step people struggle with the most to factor out the (X+4) but when you do it looks like this: \[(X+4)(X^2-4)\]

OpenStudy (anonymous):

Do you see how i got that?

OpenStudy (anonymous):

Yes. You took what they had in common and the outside of what they had in common. If that made sense. It made sense to me

OpenStudy (anonymous):

good now is there anything else to factor in that?

OpenStudy (anonymous):

Not that I see?

OpenStudy (anonymous):

what about X^2-4

OpenStudy (anonymous):

Oh yeah. It would be (x-2)(x(-2)

OpenStudy (anonymous):

Nice! so now we have: \[f(X)=(X+4)(X+2)(X-2)\] Can you tell what the zeroes are now?

OpenStudy (anonymous):

-4, -2 and 2??

OpenStudy (anonymous):

:) good job

OpenStudy (anonymous):

So the terms "roots" and "zeros" mean the same thing?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh! Thank you!

OpenStudy (anonymous):

they have similar meaning. the roots make the polynomial zero

OpenStudy (anonymous):

they get trickier when they can't be factored; luckily in this scenario wasn't the case

OpenStudy (anonymous):

Okay, that makes sense now.

OpenStudy (anonymous):

synthetic division is something i probably couldn't try to type here, that's another way of checking the roots

OpenStudy (anonymous):

I for some stupid reason can do synthetic division in my head. Just can't factor to save my life

OpenStudy (anonymous):

you're better than me i have to write it down. check with synthetic division -4, -2 and 2 into that polynomial see if you get 0

OpenStudy (anonymous):

I'm an all or nothing kind of person. XD